OpenStudy (marcelie):

help please find the area of the region that lies inside the first curve and outside the second curve

OpenStudy (marcelie):

\[r=3\cos (\theta)\] \[r=1+\cos(0)\]

OpenStudy (marcelie):

so how do i know which area is it ?

OpenStudy (marcelie):

@zepdrix

OpenStudy (3mar):

Do you mean that? \[r=1+\cos(0)\] or that? \[r=1+\cos(\theta)\]

OpenStudy (marcelie):

there are two functions :O

OpenStudy (3mar):

I mean the second one?

OpenStudy (3mar):

cos(o)=1 and 1 + cos(o)=2 so you mean r=2??

OpenStudy (3mar):

@marcelie Do you follow?

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle r_1(\theta)=3\cos(\theta) }\) \(\color{black}{\displaystyle r_2(\theta)=\cos(\theta)+1 }\) The area \(\color{black}{\displaystyle \int_{\pi/3}^{-\pi/3}\frac{1}{2}\left\{\left[3\cos(\theta)\right]^2-\left[\cos(\theta)+1\right]^2\right\}d\theta }\)

OpenStudy (solomonzelman):

(Let me recheck the setup)

OpenStudy (marcelie):

oh okay . there are two functions so i graphed it

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/y5apxfqlsc

OpenStudy (marcelie):

oh okay so which part would be shaded ?

OpenStudy (solomonzelman):

This little banana-like part on the left (the top and bottom).

OpenStudy (solomonzelman):

Oh my bad.

OpenStudy (solomonzelman):

I did inside the second curve, and outside the curve. For that my setup above would be right.

OpenStudy (solomonzelman):

For inside the first, and outside the second, you are taking the other part of the circle (not from \(\pi/3\) to \(-\pi/3\), but), from \(-\pi/3\) to \(\pi/3\). \(\color{black}{\displaystyle \int_{-\pi/3}^{\pi/3}\frac{1}{2}\left\{\left[3\cos(\theta)\right]^2-\left[\cos(\theta)+1\right]^2\right\}d\theta }\)

OpenStudy (solomonzelman):

Alternatively, if you want to note the symmetry between the top and bottom parts of the area considered, you may also say \(\color{black}{\displaystyle 2\int_{0}^{\pi/3}\frac{1}{2}\left\{\left[3\cos(\theta)\right]^2-\left[\cos(\theta)+1\right]^2\right\}d\theta }\)

OpenStudy (solomonzelman):

``` btw, correct to my previous comment: I did inside the second curve, and outside the first curve. For that my setup above would be right. ```

OpenStudy (solomonzelman):

More info ... http://tutorial.math.lamar.edu/Classes/CalcII/PolarArea.aspx (It suits most students for most purposes.)

OpenStudy (marcelie):

ohh.. so then its outside the curve - inside the curve ?

OpenStudy (solomonzelman):

The setup above is inside the first curve, and outside the second.

OpenStudy (solomonzelman):

Or, if you considered set theory ... \(r_1-r_2\). (All elements in \(r_1\) that are not in \(r_2\))

OpenStudy (solomonzelman):

I'm just trying to give better analogies of material you might have previously learned ... ignore if you didn't ..

OpenStudy (marcelie):

yeah it kinda seems familiar

OpenStudy (solomonzelman):

So, do you see why I set up the integral that way? (If you see either of the way - whether you like to use symmetry or not, it's fine, as long as you have one sufficient way of doing it0

OpenStudy (solomonzelman):

If you understand the setup we can proceed to integrate.

OpenStudy (solomonzelman):

(If not, you will need to let me know about any problems you have with it.)

OpenStudy (marcelie):

oh okay i think i get it but how did you get the bounds?

OpenStudy (solomonzelman):

Oh, I looked at the graph.

OpenStudy (solomonzelman):

You can click the lines at which the desired area starts or ends, and you will see one of the equations of y selected. In this equation, you will see the angle of boundary.

OpenStudy (solomonzelman):

(I suck at explaining how and where to click even in most critical moment.)

OpenStudy (marcelie):

lol its okay but is there an algebraic way to solve it ? in order to find the bounds

OpenStudy (solomonzelman):

You can set the curves equal to each other, and solve.

OpenStudy (solomonzelman):

\(r_1=r_2\), and then, of course, choose the solutions over the interval \([0,2\pi]\).

OpenStudy (marcelie):

i got theta = pi/3

OpenStudy (solomonzelman):

And another solution is -pi/3.

OpenStudy (solomonzelman):

Your solution is in a form \(\theta =2\pi k\pm\frac{2}{3}\pi\) ...

OpenStudy (solomonzelman):

In any case, we know where the boundaries are coming from, yes?

OpenStudy (marcelie):

oh how come ? theres another 5pi/3 right ?

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

Over \([0,2\pi]\), only the aforementioned two solutions.

OpenStudy (solomonzelman):

If we can go on with integration, let me know, if not I am willing to address more about the boundaries ... (Not rushing, just asking what you would like to do now)

OpenStudy (solomonzelman):

I guess let's integrate?

OpenStudy (solomonzelman):

I actually prefer to use the symmetry to eliminate the 1/2. \(\color{black}{\displaystyle 2\int_{0}^{\pi/3}\frac{1}{2}\left\{\left[3\cos(\theta)\right]^2-\left[\cos(\theta)+1\right]^2\right\}d\theta }\) \(\color{black}{\displaystyle \int_{0}^{\pi/3}\left\{8\cos^2(\theta)-2\cos\theta-1\right\}d\theta }\) \(\color{black}{\displaystyle \int_{0}^{\pi/3}\left\{8\left(\frac{1}{2}\cos(2\theta)+\frac{1}{2}\right)-2\cos\theta-1\right\}d\theta }\) \(\color{black}{\displaystyle \int_{0}^{\pi/3}\left\{4\cos(2\theta)-2\cos\theta+3\right\}d\theta }\)

OpenStudy (solomonzelman):

and from there simple integration, etc, although I would normally just use a calculator for this if I had to perform a quick calculation.

OpenStudy (solomonzelman):

Anyway, if any questions, ask. (I'll do my best to address them if I'm online)

OpenStudy (marcelie):

oh okay it make sense now :) tysm

OpenStudy (solomonzelman):

yw